OpenStudy (anonymous):

Radius of Convergence of power series? Σ [(2x+1)^n]/[(n+1)2^2n]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ }^{ \infty } ~ \frac{(2x+1)^n}{(n+1)(2)^{2n}}}\) like this?

OpenStudy (solomonzelman):

use the ratio test

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \lim_{n\rightarrow 0} \left|\left(A_{n+1}\right)\div\left(A_{n}\right)\right|}\)

OpenStudy (anonymous):

Like that

OpenStudy (anonymous):

Just beed help working it out I think I know the answer but not sure

OpenStudy (solomonzelman):

oh, the limit should be limit n-> infitnty

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ }^{ \infty } ~ \frac{(2x+1)^n}{(n+1)(2)^{2n}}}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \lim_{n \rightarrow \infty} \left|\frac{(2x+1)^{n+1}}{(n+1+1)(2)^{2n+1}}\times \frac{(n+1)(2)^{2n}}{(2x+1)^{n}}\right|}\) \(\large\color{black}{ \displaystyle \lim_{n \rightarrow \infty} \left|\frac{(2x+1)^{n+1}}{(n+2)(2)^{2n+1}}\times \frac{(n+1)(2)^{2n}}{(2x+1)^{n}}\right|}\) \(\large\color{black}{ \displaystyle \lim_{n \rightarrow \infty} \left|\frac{(2x+1)}{(n+2)(2)^{2n+1}}\times \frac{(n+1)(2)^{2n}}{1}\right|}\) \(\large\color{black}{ \displaystyle \lim_{n \rightarrow \infty} \left|\frac{(2x+1)}{2(n+2)}\times \frac{(n+1)}{1}\right|}\) \(\large\color{black}{ \displaystyle \left|2x+1\right| \cdot \lim_{n \rightarrow \infty} \left|\frac{n+1}{2(n+2)}\right|}\) \(\large\color{black}{ \displaystyle \frac{1}{2}\left|2x+1\right| \cdot \lim_{n \rightarrow \infty} \frac{n+1}{n+2}}\) \(\large\color{black}{ \displaystyle \frac{1}{2}\left|2x+1\right| \cdot 1}\)

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \displaystyle {\rm L}=\frac{1}{2}\left|2x+1\right| \cdot 1}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm L}=\frac{1}{2}\left|2x+1\right| }\)

OpenStudy (solomonzelman):

for series to converge \(\large\color{black}{ \displaystyle {\rm L}<1 }\) so... \(\large\color{black}{ \displaystyle \frac{1}{2}\left|2x+1\right|<1 }\) \(\large\color{black}{ \displaystyle \left|2x+1\right|<2 }\) \(\large\color{black}{ \displaystyle -2<2x+1<2 }\) \(\large\color{black}{ \displaystyle -3<2x<1 }\) \(\large\color{black}{ \displaystyle -\frac{3}{2}<x<\frac{1}{2} }\)

OpenStudy (solomonzelman):

the interval is on you to figure out

OpenStudy (freckles):

\[\large\color{black}{ \displaystyle \lim_{n \rightarrow \infty} \left|\frac{(2x+1)^{n+1}}{(n+1+1)(2)^{2(n+1)}}\times \frac{(n+1)(2)^{2n}}{(2x+1)^{n}}\right|} \] \[ \large\color{black}{ \displaystyle \frac{1}{4}\left|2x+1\right| \cdot 1}\]

OpenStudy (freckles):

\[\frac{1}{4}|2x+1|<1 \]

OpenStudy (anonymous):

The radius of convergence is 1/2 correct that is what I got, but my professor got radius of convergence is 2.

OpenStudy (freckles):

\[\frac{1}{4}|2x+1|<1 \\ \frac{1}{4}(2)|x+\frac{1}{2}|<1 \\ \frac{1}{2}|x+\frac{1}{2}|<1 \\ |x+\frac{1}{2}|<2 \]

OpenStudy (freckles):

I have the radius is 2 there

OpenStudy (anonymous):

Yes thanks